Fermi National Laboratory


Accelerator Update for December 1, 2004 - December 3, 2004

Wednesday December 1
The day shift began with the Antiproton Source (Pbar), Tevatron (TeV), Recycler, and MiniBooNE all conducting studies.

At 1:09 PM, an attempt to raise intensity in Main Injector (MI) caused a number of vacuum bursts and vacuum valves to close.

A MI expert began a three-hour slipstacking study at 2:09 PM.

Operations received permission to send beam to the Switchyard dump at 2:24 PM.

Operations reported the following on beamline status at 4:36 PM:

  • Tevatron will measure and smooth orbits, and set tunes and possibly chromaticities
  • SY beam is on hold due to power supply work
  • Recycler will first work on optimizing compensation between the Recycler and MI ramps and then work on tune coupling
  • MiniBooNE wants to take as much beam as possible
  • NuMI worked on getting beam through MI

At 8:44 PM, Operations reported the first of many MI trips due to vacuum bursts while trying to send beam to the SY dump.

Pbar experts completed their study at 11:14 PM. During their study, they managed to stack at a rate of 4mA per hour.

At 11:53 PM, the TeV began suffering from aborts at sector F2. Operators discovered a number of vacuum valves closed in the P3 line; the P3 line transfers beam from MI to SY through F sector in the TeV. Operators opened all but one of the valves from the service building.

Thursday December 2
The midnight shift began with Pbar conducting reverse proton studies, with TeV conducting ramp studies, and Recycler conducting studies, and with MiniBooNE taking beam.

At 2:37 AM, Pbar's lithium lens tripped off. It reset okay.

Pbar experts reported at 5:44 AM that they were having problems with their reverse proton study due to large beam oscillations in their transfer lines.

Operators began preparing Pbar for a tunnel access at 8:43 AM.

NuMI experts began a study period at 8:45 AM.

The Pbar access ended and they resumed stacking at 11:31 AM.

Operators prepared TeV F sector for an access at 12:36 PM.

Preaccelerator experts accessed the H- Source at 12:42 PM.

Cryo system technicians began work at 1:03 PM on wet engines at TeV sector B3 and C1.

Operations prepared Booster for an access at 1:09 PM. They came out of access an hour and a half later.

Vacuum techs reported at 2:49 PM that the TeV sector F19 valve had resumed working.

Operators reported that they had interlocked NuMI at 3:28 PM.

Operations resumed sending beam to MI, Pbar, and MiniBooNE at 3:38 PM.

Operations began sending beam to SY at 4 PM.

TeV came out of access and at 5:24 PM completed its first 980 GeV ramp. At 5:37 PM, they went into a dry squeeze. Operations established beam in the TeV at 6:06 PM.

A MI expert completed her coalescing study at 7:48 PM and a NuMI expert finished his ramp study.

A Recycler expert began a study at 9:36 PM.

Friday December 3
The midnight shift began with Pbar stacking, with the TeV, Recycler, Pbar, and SY conducting studies, and with MiniBooNE off.

Pbar stacking studies ended at 1:08 AM.

SY studies ended at 2 AM.

Recycler studies ended at 2:41 AM.

TeV experts reported at 6:08 AM that they had measured chromaticities up the ramp, but couldn't decouple the particle central orbits.

Operators prepared MiniBooNE for an access at 6:30 AM.

Plans for the Weekend
NuMI will attempt to take beam today, 12/3/04. The plan is to continue with beam startup.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/3/2004   email Fermilab